# Do LLMs have Consistent Values?

### v1

@@ 3.2 Data Analysis
When comparing to human data, we used the study in [11]. The data is from 49 cultural groups. The total number of participants was 53,472, the mean age was 34.2, (SD = 15.8), with 59% females. Their data is stored at the Open Science Framework and is available here.
@@ 3.2.1 Value Rankings
Although there is variability between individuals in their prioritization of values, there are values that tend to be ranked as more important than others across cultures and samples. Those suggest there are underlying principles that give rise to value hierarchies. The similarity in value importance across cultures is referred to as the universal value hierarchy [8, 11]. Our first question for analysis was whether this hierarchy is also reflected in LLM data. Namely, do LLMs tend to rank the same values as high or low as human subjects do.
To obtain LLM rankings for a given set of LLM answers, we assigned a score $v_{i}$ to value $i$, where $v_{i}$ was the average score given to the three items measuring this value by the LLM (ie the average of $X_{i,\cdot,\cdot}$). From this score, we subtracted the average score given to all value items within the conversation, thus centering the data. Centring is the recommended practice in value research [6, 3], and allows comparison to human samples. We then sorted these $v_{i}$ and ranked accordingly. Finally, we calculated the Spearman’s Rank Correlation ($\rho$) between this ranking and the known human ranking [11]. We also checked for significant differences between different datasets (e.g., GPT 4 with temperatures $0.0$ and $0.7$) using the Wilcoxon Signed-Rank Test.
3.2.2 Consistency Within Values
The simplest form of consistency in questionnaires is between questions that are conceptually intended to address one concept. As mentioned above, in our data there are three different questions per value. Individuals endorsing a value are likely to endorse all relevant items.
To assess this consistency here, we calculated Cronbach’s Alpha ($\alpha$) to assess how coherently LLMs express values across related items within one conversation. Cronbach’s Alpha measures the internal consistency of a scale, ranging between 0 and 1. For each value $i$ we let $\sigma^{2}_{i}$ be the variance of all $X_{i,\cdot,\cdot}$. Also let $\sigma^{2}$ denote the overall variance in $X$. Then the Cronbach alpha is:
$\alpha=\frac{V}{V-1}\left(1-\frac{\sum_{i=1}^{V}\sigma^{2}_{i}}{\sigma^{2}}\right)$ (1)
Thus, large values indicate that scores tend to be consistent within each value. To compare results for different temperatures and language models, we used paired samples t-tests to compare the average alpha across values within a dataset.
@@ 3.2.3 Correlations Between Values
The key focus of our work is correlation between values. Namely, the question of whether choice of value $i$ is correlated with that of value $j$. In humans there is a robust correlation structure where certain values are more strongly correlated than others. A standard way to represent this structure is via Multidimensional Scaling (MDS) [41], calculated as follows.
First, the matrix $C\in R^{19\times 19}$ of empirical correlation coefficients is formed. Next each of the values is embedded into $R^{2}$ via MDS, such that distances in $R^{2}$ best approximate the correlations. For human data, this results in an approximately circular embedding, as shown in [11, 9, 42]. Here we performed this analysis on the LLM data. To compare the resulting dataset to the human samples, we need to normalize for the degrees of freedom of rotation and translation. This is done via Procrustes Analysis between the human and LLM embeddings. The resulting embeddings were plotted. Then, we computed the sum of squared differences between the procrusted MDS locations of each value to the human benchmark. Larger differences indicate stronger divergence from the human samples.
Values Human Data GPT 4 Gemini Pro
Rank Mean Rank Mean Rank Mean
Benevolence-Care 1 0.79 3 0.82 5 0.90
Benevolence-Dependability 2 0.72 4 0.80 3 1.00
Self-direction Action 3 0.60 6 0.77 2 1.09
Self-direction Thought 4 0.58 5 0.78 4 0.94
Universalism-Concern 5 0.50 2 0.86 6 0.75
Universalism-Tolerance 6 0.37 1 0.97 1 1.17
Security Societal 7 0.32 11 0.16 15 -0.30
Security Personal 8 0.28 8 0.43 9 0.14
Hedonism 9 0.23 9 0.21 7 0.45
Achievement 10 0.08 14 0.04 8 0.24
Face 11 0.05 13 0.09 11 0.05
Universalism-Nature 12 -0.10 7 0.48 10 0.11
Stimulation 13 -0.11 15 -0.27 16 -0.57
Conformity-Interpersonal 14 -0.16 16 -0.55 14 -0.23
Humility 15 -0.20 12 0.15 13 -0.04
Conformity-Rules 16 -0.26 10 0.19 12 -0.02
Tradition 17 -0.72 17 -0.98 17 -1.16
Power Resources 18 -1.33 18 -2.08 18 -2.20
Power Dominance 19 -1.40 19 -2.25 19 -2.35
Table 1: Comparative analysis of the relative importance of 19 values according to LLM responses and human data. LLM resopnses are at temperature zero, and are pooled across prompts.
@@ 4 Results
Values Rankings:
Humans typically prioritize certain values over others. In this section, we analyze the LLM responses based on these rankings, as discussed in Section 3.2.1. Table 1 presents the pooled value rankings of humans and LLMs across all prompting methods. [^2] In Table 3 in the appendix we provide results for specific prompting methods (Names and Value Anchor), which show even better agreement with human rankings. The table reveals that LLM rankings generally align with human rankings in terms of which values are considered high or low. Specifically, values such as universalism, self-direction, and benevolence received high rankings, while values like tradition and power were consistently deemed less important.
To obtain a quantitative analysis of alignment of model-generated value hierarchies with human perspectives, we employed Spearman rank correlations to compare LLM and human rankings for different prompting approaches and models. These correlations are shown in Figure 2. The Basic prompt for GPT 4 shows only low correlations ($\rho$= 0.10 for $0.0$ and $\rho$=0.29 for $0.7$). This indicates that naive prompting of LLMs may be a sub-optimal approach to replicating human populations (as also discussed in [1]). In contrast, all other prompts show predominantly high associations for both GPT 4 and Gemini Pro ($\rho$ $>$ 0.77), with low variability between prompting methods. Wilcoxon Signed-Rank Tests on the pooled datasets revealed no statistically significant differences (W = 84.00, p = 0.679) between the language models GPT-4 (Pooled 00) and Gemini Pro (Pooled 00) within the same ’temperature $0.0$’ condition.
Figure 2: A heatmap of Spearman rank correlation between benchmark value hierarchies and dataset rankings for GPT 4 and Gemini Pro across temperature conditions.
Consistency Within Values:
Following Section 3.2.2, Cronbach’s $\alpha$ was used to assess the internal consistency of the value scale. For Basic Prompt $0.0$, across various models, a notable percentage of items exhibited uniformity in responses across different iterations. In the Gemini Pro model, no variability at all was found in 87.72% of the items, and negligible variability was found in 12.28% of the items. In GPT 4 only one item displayed no variability. However, the internal consistency, as indicated by Cronbach’s $\alpha$, was notably low ($\alpha$ = 0.14). Furthermore, within the Basic Prompt $0.7$, a similar trend of exceedingly low Cronbach’s $\alpha$ values was observed across various models: GPT 4 (Mean = 0.00, SD = 0.11) and Gemini Pro (Mean = -0.01, SD = 0.09).
The picture was different when using other prompts. In Figure 3, the mean and standard deviation scores of Cronbach’s $\alpha$ are visually depicted across the other models and the pooled dataset for the ’temperature $0.0$’ condition. The figure also includes the equivalent scores of a human sample, calculated as the average of the Cronbach’s $\alpha$ scores across 49 countries [11]. All of Gemini Pro’s scores in the ’temperature $0.0$’ condition were above the acceptable reliability threshold ($\alpha$ $>$ 0.60) and exhibited slightly higher scores across all prompts. In contrast, GPT 4 scores on the Demographic and Names prompts were relatively low (Mean = 0.54, SD = 0.13 and Mean = 0.45, SD = 0.25 respectively).
Figure 3: Comparing mean Cronbach’s $\alpha$ scores and standard deviations of LLMs in temperature zero across different categories with human data [11].
Increasing the temperature to $0.7$ had a negative impact on the internal consistency of both language models across prompts (see Figure 6 in the appendix). Paired-sample t-tests confirmed the significance of thiseffect for the Gemini Pro model ( t(1) = 15.97, p $<$ .001). but not for GPT 4 (t(1) = 1.00, p = .331). These results suggest that in some models elevated temperatures can compromise the coherence and consistency of language model responses.
@@
The MDS analysis (see Section 3.2.3) maps all values into $R^{2}$ in a way that reflects their correlations. Here, we conduct MDS analyses for both human responses and LLM output, and then compare the results. The analyses were performed separately for each prompt, temperature, and model. Figure 4 illustrates a comparison between human MDS and Gemini Pro at temperature $0.0$, for the Value Anchor and Names prompts, respectively. Notably, the disparities between Gemini Pro and GPT-4 for each prompt are minimal. GPT-4 plots and other Gemini Pro prompts are featured in the appendix as Figure 7 and Figure 8.
First, it can be seen that among humans, the values are organized in a circle in the theoretically expected order. These results were often identified over the years, and interpreted as resulting from the aspiration of individuals to maintain personal consistency in their motivations [6]. Second, it can be seen that the MDS configuration resulting from the Value Anchor prompt more closely follows this order than the MDS resulting from the Names prompt. We further quantitatively compared the configurations by taking the mean squared difference between any pair of human and prompting method MDS matrices (i.e., matrices in $\mathbb{R}^{19\times 2}$). See Table 2. It can be seen that the Value Anchor prompt demonstrated a better fit to human values than the other prompting methods.
@@
Figure 4: Comparison of Procrustes Analysis results between human data [11] and Gemini Pro for Value Anchor and Names prompts, for ’temperature $0.0$. The sum of squared differences, which measures the fit to human data, is 0.11 for the Value Anchor and 0.71 for the Names, indicating a better fit for the Value Anchor. For acronyms, refer to Section C.
Value Anchor Demographic Generated Persona Names
@@
00 0.11 0.42 0.39 0.71
07 0.11 0.75 0.28 0.57
@@ 5 Discussion
In the current study we analyzed the values exhibited in LLM responses. We used three metrics to estimate the quality of the LLM responses against human responses: value ranking, internal consistency, and value correlations. Our results highlighted the importance of the prompting mechanism. Using the Basic prompt (namely, just providing a questionnaire with no further instructions), the LLM was likely to either generate negligible variance across generated personas, or generate internally inconsistent outputs (respond differently to questions about the same value). These results suggest that LLMs cannot be treated as ’individuals’ holding a coherent set of value priorities.
@@
Prompts that endow the LLM with a “personality” improved the consistency of each specific value profile, to varying degrees. The value hierarchy was consistently found across prompts, indicating that at the mean level, LLMs can simulate value rankings of human populations. More variability between conditions was found in the internal consistency metric, with the Values Anchor prompt providing the best consistency, and the Names prompt a low consistency that would be considered unacceptable in a human sample. Finally, the Value Anchor prompt also best modeled the inter-value correlations. This is arguably the most important metric since it allows analyzing consistency across values, within a single session. These results suggest that LLMs, applying suitable prompts, can produce a ’population’ of individuals, each reporting a different, but coherent set of value priorities.
@@
One fascinating question is where the LLM learns to produce such clear profiles of values. These profiles may be implicitly learned during pre-training. Indeed, past studies indicated that values can be identified in texts, such as newspaper articles and social media. However, these values did not necessary follow the theoretical value inter-relations identified here [43, 44, 45]. Individuals who value competing values may experience stress and indecision when faced with a dilemma, resulting in gradual change in values toward a more coherent form [46, 42]. In contrast, text may very well present both sides of a dilemma and thus retain inconsistencies. Such inconsistencies were not identified because past studies relied mostly on lexical approaches [43, 44, 45]. LLMs, taking context into account, may be more likely to identify value inter-relations correctly. LLMs may also have learned to produce value profiles in the process of fine-tuning or RLHF [47]. Future research can try to distinguish these two sources of learning using careful analysis of training sources as well as evaluation of different checkpoints in the training process.
@@
Past research into human personas sought ways to estimate the ability of the LLM to maintain a consistent persona across a conversation [34]. We establish that the unique qualities of human values, and the ample empirical knowledge collected about them, allow their use as such a method [3, 4, 48]. We suggest that known behavioral correlations in humans can be applied to assess the consistency of LLM personas. Here we focused on evaluations via a questionnaire, but one could envision more elaborate evaluations that rely on other features of human personalities.
@@
The procedures and data produced here may have important contributions for psychological research. Investigators interested in human behavior can apply these procedures to produce datasets that simulate human samples. Future research can investigate their possible use to replicate known findings (e.g., age differences in values) or pretest novel hypotheses (e.g. associations between values and specific behaviors).
@@
The question of values in LLMs is of course of philosophical and societal importance. Our results show that on average, these values largely reproduce international value rankings. However, small variations in value importance may have implications at the societal and individual level (e.g. gender roles [49]; entrepreneurship [50]; prosocial behavior [51]; and antisocial behavior [52]). Future work should consider the influence of these values on LLM responses, as well as on the individuals interacting with them.
@@
The current study focused on a limited number of contexts (e.g., five prompts, two temperatures, and two models). Importantly, we found commonalities across the various contexts, beside the differences. Future studies can use these results to understand what other contexts should be investigated, to possibly further enhance the quality of the output. Another limitation is the restriction to one questionnaire (PVQ-RR) corresponding a specific value system. It will be interesting to explore other forms of probing values.

### v3

@@ 3.2 Data Analysis
When comparing to human data, we used the study in Schwartz and Cieciuch (2022). The data is from 49 cultural groups. The total number of participants was 53,472, the mean age was 34.2, (SD = 15.8), with 59% females. Their data is stored at the Open Science Framework and is available here.
@@ 3.2.1 Value Rankings
Although there is variability between individuals in their prioritization of values, there are values that tend to be ranked as more important than others across cultures and samples. Those suggest there are underlying principles that give rise to value hierarchies. The similarity in value importance across cultures is referred to as the universal value hierarchy (Schwartz and Bardi, 2001, Schwartz and Cieciuch, 2022). Our first question for analysis was whether this hierarchy is also reflected in LLM data. Namely, do LLMs tend to rank the same values as high or low as human subjects do.
To obtain LLM rankings for a given set of LLM answers, we assigned a score $v_{i}$ to value $i$, where $v_{i}$ was the average score given to the three items measuring this value by the LLM (i.e. the average of $X_{i,\cdot,\cdot}$). From this score, we subtracted the average score given to all value items within the conversation, thus centering the data. Centring is the recommended practice in value research (Schwartz, 1992, Sagiv and Schwartz, 2022), and allows comparison to human samples. We then sorted these $v_{i}$ and ranked accordingly. Finally, we calculated the Spearman’s Rank Correlation ($\rho$) between this ranking and the known human ranking (Schwartz and Cieciuch, 2022). We note that this analysis does not consider correlations between answers given in the same session, and thus it may be viewed as analyzing the first-order statistics of the responses.
@@ 3.2.2 Correlations Between Values
A key focus of our work is correlation between values. Namely, the question of whether choice of value $i$ is correlated with that of value $j$. In humans there is a robust correlation structure where certain values are more strongly correlated than others. A standard way to represent this structure is via Multidimensional Scaling (MDS) (Borg et al., 2018), calculated as follows.
First, the matrix $C\in R^{19\times 19}$ of empirical correlation coefficients is formed. Next each of the values is embedded into $R^{2}$ via MDS, such that distances in $R^{2}$ best approximate the correlations. For human data, this results in an approximately circular embedding, as shown in (Schwartz and Cieciuch, 2022, Skimina et al., 2021a, Daniel and Benish-Weisman, 2019). Here we performed this analysis on the LLM data. To compare the resulting dataset to the human samples, we need to normalize for the degrees of freedom of rotation and translation. This is done via Procrustes Analysis between the human and LLM embeddings. The resulting embeddings were plotted. Then, we computed the sum of squared differences between the procrusted MDS locations of each value to the human benchmark. Larger differences indicate stronger divergence from the human samples.
@@ 4 Results
The above analyses were performed for all models and prompting strategies. We checked that model responses only contained scores for the questions in the questionnaires, and that they could therefore be transformed to tabular form and analyzed. This was almost always the case except for Gemma 2 27B on the Demographic prompt at temperature zero, and we therefore do not provide result for that settings.
Value Rankings:
As previously mentioned, research across samples and cultures have shown that while individual differences exist in human value priorities, there are also robust common patterns. In this section, we analyze the LLM responses and compare them to the typical ranking of human values, as discussed in Section 3.2.1.
Figure 2(a) shows the Spearman rank correlations between human rankings and those of the different models and prompting schemes. The results show high correlation levels ($>0.8$) for many prompt-model combinations. One exception is the basic prompt with GPT, which shows very low correlation. Full rankings are provided in the Appendix for several models and prompts (Table 2, Table 3 and Table 4). These reveal that values such as Benevolence that are highly ranked in humans are indeed also highly ranked by most LLMs (e.g,. ranked third and first by GPT-4 for the Value Anchor prompt with temperatures 00 and $0.7$ respectively). Conversely, values such as Power Dominance that are ranked low by humans, are ranked low by models (e.g., 19 by GPT-4 for the Value Anchor prompt). Figure 2(b) shows the scores corresponding to the Value Anchor prompt, when sorted according to human preferences. It can be seen that the models tend to agree with the human ordering on the low and high ranked values. Taken together, these results demonstrate that LLMs tend on average to align with the human ranking of values.
(a) Spearman Rank Correlation Coefficients
(b) Value Ranking
Figure 2: Left: A heatmap of Spearman rank correlation between benchmark value hierarchies and dataset rankings for GPT 4, Gemini Pro, Llama 3.1 8B and 70B instruct, and Gemma 2 9B and 27B across temperature conditions. Right: Average value scores for the Value Anchoring prompt at zero temperature. The x-axis shows values ordered according to human ranking (i.e., Power ranks lowest for humans and Benevolence ranks highest). The y-axis is the mean-centered scores the models ascribe to these values in the questionnaire, and human values in red. It can be seen that models tend to give lower scores to values that are ranked lower by humans, and higher scores to values ranked higher. The LLM scores also track the human scores (red curve) quite well.
@@
The MDS analysis (see Section 3.2.2) maps all values into $R^{2}$ in a way that reflects their correlations. Here, we conduct MDS analyses for both human responses and LLM output, and then compare the results. The analyses were performed separately for each prompt, temperature, and model. Figure 3 illustrates a comparison between human MDS and Gemini Pro at temperature $0.0$, for the Value Anchor and Names prompts, respectively. Notably, the disparities between Gemini Pro and GPT-4 for each prompt are minimal. GPT-4 plots and other Gemini Pro prompts are included in the Appendix as Figure 7 and Figure 8.
First, it can be seen that among humans, the values are organized in a circle in the theoretically expected order. These results were often identified over the years, and interpreted as resulting from the aspiration of individuals to maintain personal consistency in their motivations (Schwartz, 1992). Second, it can be seen that the MDS configuration resulting from the Value Anchor prompt more closely follows this order than the MDS resulting from the Names prompt. We further quantitatively compared the configurations in Table 1, by taking the mean squared difference between any pair of human and prompting method MDS matrices (i.e., matrices in $\mathbb{R}^{19\times 2}$). It can be seen that the Value Anchor prompt demonstrated a better fit to human values than the other prompting methods.
@@
Figure 3: Comparison of Procrustes Analysis results between human data (Schwartz and Cieciuch, 2022) and Gemini Pro for Value Anchor and Names prompts, for temperature $0.0$. The sum of squared differences, which measures the fit to human data, is 0.11 for the Value Anchor and 0.71 for the Names, indicating a better fit for the Value Anchor. For acronyms, refer to Section C.
Basic Value Anchor Demographic Persona Names
@@
00 0.87 0.11 0.42 0.39 0.71
07 0.69 0.11 0.75 0.28 0.57
Llama 3.1 8B
00 0.80 0.18 0.47 0.58 0.60
07 0.57 0.16 0.47 0.58 0.57
Llama 3.1 70B
00 0.61 0.10 0.29 0.37 0.45
07 0.44 0.10 0.22 0.40 0.44
Gemma 2 9B
00 0.42 0.10 0.19 0.39 0.23
07 0.82 0.11 0.16 0.32 0.12
Gemma 2 27B
00 NA 0.16 NA 0.31 0.23
07 0.64 0.17 0.15 0.25 0.19
Table 1: Sum of squared difference between the MDS embeddings of humans and LLM. Gemma 2 27B did not produce parseable results for the Demographic prompt, and for Gemma 27B at temperature 0, some values had zero-variance, thus precluding computation of correlation coefficients. All Llamma models are Instruct.
4.1 Understanding Value Anchoring
The results above show that value anchoring generates a correlation structure between values that is in better agreement with that of humans. We next set out to understand why this is the case. As we shall show, anchoring on a value not only increases the score the model assigns to the value, but it only changes the way other values are scored. Specifically, values that are close to the anchored value (on the value circle, Figure 1) tend to receive high scores. On the other hand, values that are far from the anchor receive consistently lower scores. This in turn has the effect of increasing correlation between values in the anchoring setting.
To show the above, we produce an “anchored score curve” as follows. First, we order the 19 anchoring values according to their order on the value circle in Figure 1. Note that in this order, values 19 and 1 will actually be close in the circle. Then, for each set of responses of an LLM to a Value Anchor prompt, we shift the anchored value to zero. We then average all these shifted curves. Results for all models are shown in Figure 4, along with a sine function which shows a good fit to these curves. The behavior of all models is quite consistent, except for Gemma-2-9B.
As expected, the anchored value receives the highest score. However, what is more interesting is that the values close to it tend to get similarly high, and values farther away (e.g., 180 degrees apart) receive the lower values. This means that the anchored model scores values in a way that is consistent with its anchoring. This in turn implies that neighboring values will tend to be correlated, thus explaining why Value Anchoring better captures human correlation patterns.
Figure 4: Analysis of scores after value anchoring. The plot shows the average of the score values after shifting to the anchored value. It can be seen that the anchored value receives the highest score, as expected. More surprisingly, neighboring values receive similarly high values, whereas more distant values receive lower values.
@@ 5 Discussion
In the current study we analyzed the values exhibited in LLM responses. We used two metrics to estimate the quality of the LLM responses against human responses: value ranking, and value correlations. Our results highlighted the importance of the prompting mechanism. Using the Basic prompt (namely, just providing a questionnaire with no further instructions), the LLM was likely to either generate negligible variance across generated personas, or generate internally inconsistent outputs (respond differently to questions about the same value). These results suggest that LLMs cannot be treated as ’individuals’ holding a coherent set of value priorities. In contrast, our results indicated high consistency across model types, including commercial and open LLMs.
@@
Prompts that endow the LLM with a “personality” improved the consistency of each specific value profile, to varying degrees. The value hierarchy was consistently found across prompts, indicating that at the mean level, LLMs can simulate value rankings of human populations. More variability between conditions was found in the measure of inter-value correlations. This is arguably the most important metric since it allows analyzing consistency across values, within a single session. We found best consistency in the Values Anchor prompt. These results suggest that LLMs, applying suitable prompts, can produce a ’population’ of individuals, each reporting a different, but coherent set of value priorities.
@@
It is important to note that in neither of the prompts did the LLMs receive instructions for answering about all values. The results suggest that LLM is not only instruction-following, but uses the instructions as a context that guides consistent answers with relations to a variety of values. One fascinating question is where the LLM learns to produce such clear profiles of values. These profiles may be implicitly learned during pre-training. Indeed, past studies indicated that values can be identified in texts, such as newspaper articles and social media. However, these values did not necessary follow the theoretical value inter-relations identified here (Bardi et al., 2008, Ponizovskiy et al., 2020, Kumar et al., 2018). Individuals who value competing values may experience stress and indecision when faced with a dilemma, resulting in gradual change in values toward a more coherent form (Bardi et al., 2009, Daniel and Benish-Weisman, 2019). In contrast, text may very well present both sides of a dilemma and thus retain inconsistencies. Such inconsistencies were not identified because past studies relied mostly on lexical approaches (Bardi et al., 2008, Ponizovskiy et al., 2020, Kumar et al., 2018). LLMs, taking context into account, may be more likely to identify value inter-relations correctly. LLMs may also have learned to produce value profiles in the process of fine-tuning or RLHF (Qiu et al., 2022). Future research can try to distinguish these two sources of learning using careful analysis of training sources as well as evaluation of different checkpoints in the training process.
@@
Past research into human personas sought ways to estimate the ability of the LLM to maintain a consistent persona across a conversation (Wang et al., 2024). We establish that the unique qualities of human values, and the ample empirical knowledge collected about them, allow their use as such a method (Sagiv and Schwartz, 2022, Sagiv et al., 2017, Knafo-Noam et al., 2024). We suggest that known behavioral correlations in humans can be applied to assess the consistency of LLM personas. Here we focused on evaluations via a questionnaire, but one could envision more elaborate evaluations that rely on other features of human personalities.
@@
The procedures and data produced here may have important contributions for psychological research. Investigators interested in human behavior can apply these procedures to produce datasets that simulate human samples. Future research can investigate their possible use to replicate known findings (e.g., age differences in values) or pretest novel hypotheses (e.g. associations between values and specific behaviors). The use of both commercial and open LLMs increases the reproducibilty of the results.
@@
The question of values in LLMs is of course of philosophical and societal importance. Our results show that on average, these values largely reproduce international value rankings. However, small variations in value importance may have implications at the societal and individual level (e.g. gender roles (Lomazzi and Seddig, 2020); entrepreneurship (Woodside et al., 2020); prosocial behavior (Daniel et al., 2020); and antisocial behavior (Benish-Weisman, 2019)). Future work should consider the influence of these values on LLM responses, as well as on the individuals interacting with them.
@@
The current study focused on a limited number of contexts (e.g., five prompts, two temperatures, and six models). Importantly, we found commonalities across the various contexts, beside the differences. Future studies can use these results to understand what other contexts should be investigated, to possibly further enhance the quality of the output. Another limitation is the restriction to one questionnaire (PVQ-RR) corresponding a specific value system. It will be interesting to explore other forms of probing values.
@@
Table 3: Comparative analysis of values’ relative importance for the Llama 3.1 8B and Llama 3.1 70B datasets across temperatures.
Benchmark Llama 3.1 8B Llama 3.1 70B
1 BEC 0.79 0.74 4 0.87 3 0.62 8 0.82 4 0.80 2 0.80 2 0.87 3 0.83 5
2 BED 0.72 0.64 5 0.64 5 0.62 9 0.65 7 0.64 4 0.69 4 0.67 7 0.69 7
3 SDA 0.60 1.00 2 1.04 2 1.12 2 1.12 2 0.76 3 0.72 3 0.84 5 0.84 4
4 SDT 0.58 0.24 11 0.34 9 0.89 5 0.67 6 0.48 5 0.53 5 1.01 2 0.96 2
5 UNC 0.50 0.77 3 0.80 4 1.11 3 0.99 3 0.46 6 0.49 6 0.85 4 0.84 3
6 UNT 0.37 1.14 1 1.12 1 1.19 1 1.20 1 0.90 1 0.88 1 1.10 1 1.10 1
7 SES 0.32 0.34 9 0.47 8 0.91 4 0.81 5 -0.01 11 0.01 10 0.05 11 0.08 11
8 SEP 0.28 0.62 6 0.58 6 0.84 6 0.64 8 0.23 9 0.23 9 0.34 8 0.39 8
9 HE 0.23 -0.61 16 -0.58 16 -1.38 16 -0.50 14 -0.51 16 -0.46 16 -0.13 13 -0.31 13
10 AC 0.08 -0.37 13 -0.36 13 -0.27 13 -1.20 12 0.32 8 0.26 8 0.23 10 0.24 9
11 FAC 0.05 -0.46 14 -0.58 15 0.75 15 -0.63 16 -0.24 14 -0.35 14 -0.60 15 -0.66 15
12 UNN -0.10 0.55 7 0.53 7 0.56 11 0.48 9 0.01 10 -0.03 11 0.29 9 0.15 10
13 ST -0.11 -0.59 15 -0.53 14 -0.67 14 -0.34 13 -0.15 13 -0.12 13 -0.03 12 -0.08 12
14 COI -0.16 -0.00 12 -0.14 12 0.61 10 0.29 10 -0.30 15 -0.37 15 -0.89 17 -0.74 16
15 HUM -0.20 0.44 8 0.29 10 0.69 7 0.47 10 0.35 7 0.29 7 0.74 6 0.81 6
16 COR -0.26 0.29 10 0.22 11 0.56 12 0.03 11 -0.14 12 -0.11 12 -0.35 14 -0.36 14
17 TR -0.72 -1.28 17 -1.28 17 -2.01 18 -1.53 17 -0.68 17 -0.66 17 -0.82 16 -0.82 17
18 POR -1.33 -1.35 18 -1.29 18 -1.77 17 -1.55 18 -0.89 18 -0.88 18 -1.46 18 -1.46 18
19 POD -1.40 -2.12 19 -2.11 19 -2.88 19 -2.56 19 -1.86 19 -1.81 19 -2.39 19 -2.36 19
Table 4: Comparative analysis of values’ relative importance for the Gemma 2 9B and Gemma 2 27B datasets across temperatures.
Benchmark Gemma 2 9B Gemma 2 27B
1 BEC 0.79 1.14 2 1.22 2 1.90 2 1.88 2 1.57 1 1.57 1 1.53 1 1.50 1
2 BED 0.72 0.73 5 0.78 5 1.29 4 1.24 4 1.16 2 1.17 2 1.29 3 1.25 3
3 SDA 0.60 1.55 1 1.58 1 1.77 3 1.77 3 0.84 4 0.84 4 1.50 2 1.46 2
4 SDT 0.58 0.90 4 0.89 3 1.93 1 1.91 1 0.30 8 0.27 8 1.20 4 1.18 4
5 UNC 0.50 0.67 6 0.65 6 1.03 6 1.05 5 0.83 5 0.81 5 1.02 6 1.04 5
6 UNT 0.37 0.94 3 0.84 4 1.03 5 1.02 6 0.92 3 0.89 3 1.02 5 1.02 6
7 SES 0.32 -0.01 10 0.00 10 0.75 8 0.76 8 0.11 10 0.12 10 0.32 10 0.33 9
8 SEP 0.28 0.08 9 0.05 9 0.15 9 0.12 9 0.07 11 0.07 11 0.11 11 0.07 11
9 HE 0.23 -0.33 15 -0.32 15 -0.54 13 -0.55 13 -0.23 13 -0.23 13 -0.55 13 -0.49 13
10 AC 0.08 -0.22 13 -0.22 13 -0.50 12 -0.50 12 -0.19 12 -0.21 12 -0.34 12 -0.32 12
11 FAC 0.05 -1.01 16 -1.05 16 -1.45 17 -1.45 17 -0.84 16 -0.81 16 -1.24 17 -1.24 17
12 UNN -0.10 0.21 8 0.20 8 -0.03 10 0.01 10 0.56 6 0.45 6 0.55 7 0.61 7
13 ST -0.11 -0.27 14 -0.24 14 -0.72 14 -0.77 14 -0.49 15 -0.42 15 -0.72 14 -0.77 15
14 COI -0.16 -0.21 12 -0.20 12 -0.94 15 -0.93 15 -0.33 14 -0.31 14 -1.00 16 -1.01 16
15 HUM -0.20 -0.14 11 -0.18 11 -0.40 11 -0.41 11 0.13 9 0.13 9 0.47 8 0.47 8
16 COR -0.26 0.55 7 0.63 7 0.84 7 0.85 7 0.30 7 0.29 7 0.34 9 0.31 10
17 TR -0.72 -1.23 17 -1.24 17 -0.99 16 -0.96 16 -0.99 17 -0.96 17 -0.80 15 -0.76 14
18 POR -1.33 -1.63 18 -1.62 18 -2.53 18 -2.48 18 -1.41 18 -1.41 18 -2.00 18 -1.96 18
19 POD -1.40 -1.72 19 -1.77 19 -2.59 19 -2.57 19 -2.30 19 -2.26 19 -2.70 19 -2.69 19